The fastest speed ever measured for a tennis ball served by a player was 263 km/h. At this speed, how long does the tennis ball take to go from one end of the court to the other, a distance of 23.8 m?
Answer #1

Speed =263 km/h = 263×1000/60×60= 73.0556 m/sec

Distance S = 23.8 m.

Time T = distance/ speed.

T= 23.8/73.0556.

T = 0.32577 sec.